OpenStudy (anonymous):

what is the GCF of 18x^2y^4 and 24x^4y^4

OpenStudy (jdoe0001):

\(\large \begin{cases} 18x^2y^4 \\ 24x^4y^4 \end{cases}\implies \begin{cases} {\color{brown}{ 2}}\cdot {\color{brown}{ 3}}\cdot 3{\color{brown}{ x^2y^4}}\\ {\color{brown}{ 2}}\cdot 2\cdot 2\cdot {\color{brown}{ 3}}\cdot {\color{brown}{ x^2y^4}} \end{cases}\) any ideas?